Insert 2 geometric means between -32 and 4
Answer by stanbon(75887)

Insert 2 geometric means between -32 and 4
------
Progression: -32,-32x,-32x^2,-32x^3=4
--------------------------------------
Solve for "x" and "x^2)
-32x^3 = 4
-----
x^3 = -1/8
x = (-1/2)
x^2 = (1/4)
-------
Answer:: -32x = -32(-1/2) = 16 and -32x^2 = -32(1/4) = -8
------------
Cheers,
Stan H.
---------------------